Two cities n miles a part are located s inches a part on a certain map that is drawn to scale. What is the distance, in inches, on the map between two cities that are n+1 miles apart?
A n/s B(n+1)/s C s/(n+1) D s(n+1)/n E n/s(n+1

Answer by edjones(8007)

let d be the distance, in inches, on the map between two cities that are n+1 miles apart
n/s = (n+1)/d
dn=s(n+1)
d=s(n+1)/n
.
Ed